An optical grating with 5000 line 51cm gives a second order iamge. Ata diffraction angle 44.8 degree calculate the wave. Lenght of the light used
1
Expert's answer
2012-11-20T11:26:48-0500
wavelength =period*sin(angle)/order = width*sin(angle)/(lines*order) = (51 cm)*sin(44.8
deg)/(5000*2)
= 3.594 * 10^-5 m = 35940 nm

Answer: 35940 nanometers or 35.94 micrometers